### v4

# STORM: A Spatio-Temporal Factor Model Based on Dual Vector Quantized Variational Autoencoders for Financial Trading

Abstract.
In financial trading, factor models are widely used to price assets and capture excess returns from mispricing. Recently, we have witnessed the rise of variational autoencoder-based latent factor models, which learn latent factors self-adaptively. While these models focus on modeling overall market conditions, they often fail to effectively capture the temporal patterns of individual stocks. Additionally, representing multiple factors as single values simplifies the model but limits its ability to capture complex relationships and dependencies. As a result, the learned factors are of low quality and lack diversity, reducing their effectiveness and robustness across different trading periods. To address these issues, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders, named STORM, which extracts features of stocks from temporal and spatial perspectives, then fuses and aligns these features at the fine-grained and semantic level, and represents the factors as multi-dimensional embeddings. The discrete codebooks cluster similar factor embeddings, ensuring orthogonality and diversity, which helps distinguish between different factors and enables factor selection in financial trading. To show the performance of the proposed factor model, we apply it to two downstream experiments: portfolio management on two stock datasets and individual trading tasks on six specific stocks. The extensive experiments demonstrate STORM’s flexibility in adapting to downstream tasks and superior performance over baseline models. Code is available in PyTorch [^1] https://github.com/DVampire/Storm.
Factor Model, VQVAE, Portfolio Management, Algorithmic Trading
journalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3773966.3777972isbn: 979-8-4007-2292-9/2026/02ccs: Information systems Spatial-temporal systemsccs: Applied computing Electronic commerce
1. Introduction
In financial trading, factor models are fundamental tools for asset pricing and are widely used to predict asset returns. These models enhance pricing accuracy and risk management by identifying a set of key factors that explain excess returns. The well-known Fama-French three-factor model (fama1992cross) extends the Capital Asset Pricing Model (sharpe1964capital) with additional factors to capture market risk premiums. As financial theories evolve, researchers continue to discover new factors, broadening our understanding of market dynamics.
Figure 1. Overall architecture of STORM
Recently, we have witnessed the rise of latent factor models (gu2021autoencoder; duan2022factorvae; wei2022factor; wei2023hirevae), connecting the factor model with the generative model, the variational autoencoder (VAE). These VAE-based models describe high-dimensional data (prices) to low-dimensional representations (factors), and learn factors self-adaptively. Although latent factor models have demonstrated substantial success in financial trading tasks, they still face several significant issues:
  • CH1: Limited Reflection of Market Complexity. Latent factor models represent factors as single values are inherently constrained by their insufficient capacity to capture the intricate complexity and nonlinearity of financial data, rendering them vulnerable to noise and non-stationarity, which compromises their predictive accuracy and stability.
  • CH2: Factor Inefficiency. VAE-learned factors suffer from three inefficiencies: i) focusing mainly on cross-sectional factors while neglecting temporal information, ii) allowing noise in continuous latent spaces to overshadow meaningful signals, and iii) lacking independence among factors, which leads to multicollinearity and weak adaptability to varying market conditions.
  • CH3: Lack of Factor Selection. Existing latent factor models primarily focus on generating factors without adequately differentiating between them. Furthermore, they neglect the crucial process of factor selection, which is essential for identifying impactful factors, thereby limiting the model’s overall effectiveness and precision.
In order to address the challenges, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders (VQ-VAE), named STORM, with the architecture shown in Figure 1. Unlike traditional scalar-valued financial factors with clear economic interpretations, STORM learns high-dimensional latent vectors that, while less interpretable capture complexity and nonlinearity inherent in financial martket (CH1). Additionally, we develop a dual VQ-VAE architecture to capture cross-sectional and time-series features, considering both spatial and temporal perspectives [^2] We treat cross-sectional as spatial and time-series as temporal. These terms will be used interchangeably unless specified.. By integrating these features at both fine-grained and semantic levels, the model constructs more effective factors. Through diversity and orthogonality loss constraints, we ensure representation independence among learned representations (CH2). Furthermore, codebook embeddings act as cluster centers, serving as class tokens to categorize factor embeddings. This strategy provides clarity and transparency to the differentiation and selection process of factors (CH3). Specifically, our contributions are four-fold:
  • We design a dual VQ-VAE architecture that systematically constructs cross-sectional and time-series factors by capturing both spatial and temporal dependencies.
  • We leverage vector quantization techniques to improve factor embedding representation by reducing noise, enhancing diversity, and ensuring orthogonality. This approach significantly strengthens the factors’ predictive power and a more transparent factor selection process.
  • We develop STORM as a versatile all-rounder capable of supporting three tasks simultaneously, whereas other methods tend to be specialized in specific tasks.
  • Experiments on seven datasets from two U.S. stock markets, covering the stock future return prediction task (DJ30 and SP500), portfolio management (DJ30 and SP500), and algorithmic trading (five stocks), show that STORM outperforms all baselines across two prediction metrics and six standard financial metrics.
2. Related Work
2.1. Factor Model
Factor model has attracted extensive research focus from the finance community. Financial experts have carefully selected a wide range of factors from macroeconomic (sharpe1964capital), market changes (fama1992cross), individual stock fundamentals (greenblatt2010little), and other perspectives to price assets in an attempt to capture excess returns from mispricing. Recently, latent factor models have emerged that improve investment returns by studying latent factors self-adaptively. Specifically, (gu2021autoencoder) learns latent factors depend on asset characteristics, (duan2022factorvae) extracts cross-sectional factors by predicting future returns, and (wei2023hirevae) learns market-augmented latent factors. However, these VAE-based latent factor models may suffer from a lack of robustness in representing low signal-to-noise ratio markets, leading to reduced reconstruction and generation capabilities. Furthermore, their effectiveness is constrained by an insufficient focus on the temporal perspective.
2.2. Financial Trading
Financial trading has attracted significant attention from both the finance and AI communities. Quantitative traders leverage mathematical models and algorithms to automatically identify trading opportunities (sun2024trademaster). Within the FinTech domain, key research areas encompass portfolio management (PM), algorithmic trading (AT), order execution (fang2021universal), market making (spooner2020robust), etc. In this work, we focus on PM and AT, as they are fundamental to developing intelligent financial decision-making systems.
PM focuses on optimizing wealth allocation across assets (li2014online). Its development has evolved from simple rule-based methods (poterba1988mean) to prediction-based approaches using machine learning (ke2017lightgbm) and deep learning (qin2017dual; zhao2022stock) for predicting returns or price movements. Despite recent integration of reinforcement learning (RL) in PM (zhang2024reinforcement), the latent factor models still predominantly rely on deep learning for market modeling and precise predictions.
AT, on the other hand, involves executing trades with algorithm-generated signals. Like PM, AT has seen a transition from rule-based and supervised learning techniques (yang2020qlib) to RL-driven approaches. In particular, RL-based methods (sun2022deepscalper) have gained prominence due to their ability to navigate complex sequential decision-making problems, further enhancing trade execution efficiency and adaptability in dynamic market environments.
2.3. Vector Quantized Variational Autoencoder
Vector Quantized Variational Autoencoder (VQ-VAE) has emerged as a fundamental framework in deep learning research, establishing methodological significance across core domains: representation learning, generative modeling, and time series analysis. Initially introduced by van2017neural(van2017neural), VQ-VAE has showcased its capabilities in processing high-dimensional data across diverse modalities, including images (razavi2019generating), audio (zhang2024codebook), and video (yan2021videogpt), with a strong emphasis on learning discrete latent representations. Particularly noteworthy is the model’s demonstrated proficiency in capturing temporal dependencies, where recent studies like (talukder2024totem) have validated its superior performance in modeling sequential patterns compared to conventional continuous latent space approaches, thereby advancing the development of robust generative architectures specifically optimized for temporal dynamics analysis.
3. Preliminaries
3.1. Factor Model
The factor model (e.g., Arbitrage Pricing Theory (ross1976arbitrage)) is defined to study the relationship between expected asset returns and factor exposures, focusing on the cross-sectional differences in expected asset returns.
(1) $\mathbb{E}[R_{i}^{e}]=\alpha_{i}+\beta_{i}\lambda\text{.}$
Here, $\mathbb{E}[\cdot]$ denotes the expectation operator, and $\mathbb{E}[R_{i}^{e}]$ represents the expected excess return of asset $i$. $\beta_{i}$ is factor exposure (or factor loading) of asset $i$, and $\lambda$ is factor expected return (or factor risk premium). Lastly, $\alpha_{i}$ captures pricing error, reflecting any deviations from the model’s predicted returns.
According to (duan2022factorvae), the factor model is formulated from a spatial perspective, primarily examining cross-sectional variations in expected returns across different assets, rather than capturing the temporal evolution of individual asset returns. Therefore, we expand it from the temporal perspective and introduce the general functional form of the latent factor model:
(2) $y_{i,t}=\alpha_{i,t}+\sum_{k=1}^{K}\bm{\beta}_{i,t}^{k}\mathbf{z}_{i,t}^{k}+\epsilon_{i,t}\text{,}$
where $y_{i,t}$ denotes the future return of stock $i$ at timestep $t$, $\mathbf{z}_{i,t}$ denotes $K$ factor returns, and $\epsilon_{i,t}$ is the noise. From the spatial, Equation 2 can be expressed as $\mathbf{y}_{i}=\bm{\alpha}_{i}+\sum_{k=1}^{K}\bm{\beta}_{i}^{k}\mathbf{z}_{i}^{k}+\bm{\epsilon}_{i}$, which we regard as the cross-sectional factor model. From the temporal perspective, Equation 2 can be expressed as $\mathbf{y}_{t}=\bm{\alpha}_{t}+\sum_{k=1}^{K}\bm{\beta}_{t}^{k}\mathbf{z}_{t}^{k}+\bm{\epsilon}_{t}$, which we regard as the time-series factor model. The cross-sectional factor model focuses on the differences in returns of different assets at the same point in time, while the time-series factor model focuses on the changes in returns of a single asset at different time points.
3.2. Problem Formulation
To assess the effectiveness and generating ability of factors, we predict and evaluate the stock future returns. To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (wang2021deeptrader; zhang2024reinforcement; ye2020reinforcement) and algorithmic trading (sun2022deepscalper; liu2020adaptive), to showcase the model’s investment capabilities.
Observed Data. We utilize the stock’s historical price data $\mathbf{p}$ (i.e., open, high, low, and close) and technical indicators $\mathbf{d}$ as observed variables $\mathbf{x}:=[\mathbf{p},\mathbf{d}]\in\mathbb{R}^{N\times W\times D}$, within a window size of $W$. The historical price data of stock $i$ at time $t$ is denoted as $\mathbf{p}_{i,t}:=[p_{i,t}^{o},p_{i,t}^{h},p_{i,t}^{l},p_{i,t}^{c}]\in\mathbb{R}^{D_{1}}$. The technical indicators which are calculated from price and volume values, are denoted as $\mathbf{d}_{i,t}\in\mathbb{R}^{D_{2}}$. Each stock is represented by $D=D_{1}+D_{2}$ features per trading day, which we then use to predict the stock’s future returns, defined as $\mathbf{y}_{i,t+1}:=\frac{p_{i,t+1}-p_{i,t}}{p_{i,t}}$.
Downstream tasks. These primarily encompass the two most widely studied areas in quantitative finance – Portfolio management and Algorithmic trading:
  • Portfolio management task aims to construct an optimal portfolio based on the predicted returns to test the profitability of the factor model. At each trading day $t$, the factor model outputs return predictions $\hat{\mathbf{y}}_{i,t+1}$ for each stock $i$, which are then used by a portfolio allocation strategy, e.g., TopK-Drop, to generate the portfolio weight vector $\mathbf{w}_{t}=[w_{t}^{0},w_{t}^{1},\dots,w_{t}^{N-1}]\in\mathbb{R}^{N}$, where $w_{t}^{i}$ represents the proportion of capital invested in stock $i$. To ensure full capital deployment, the allocation must satisfy the budget constraint $\sum_{i=0}^{N-1}w_{t}^{i}=1$.
  • Algorithmic trading task aims to execute buy, hold, and sell actions based on predicted asset states to balance returns and risks. We formulate it as a Markov Decision Process (MDP) under reinforcement learning scenario following (chen2023mtrader; zhang2024finagent). The MDP is constructed by a 5-tuple $(\mathit{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where $\mathit{S}$ and $\mathcal{A}$ are sets of states and actions respectively, $\mathcal{T}:\mathit{S}\times\mathcal{A}\times\mathit{S}\rightarrow[0,1]$ is the state transition function, $\mathcal{R}:\mathit{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function where $\mathbb{R}$ is a continuous set of possible rewards, and $\gamma\in[0,1)$ is the discount factor. The latent factor embeddings $\mathbf{Z}$ generated by the factor model are integrated into the observation set that constitutes the state space $\mathcal{S}$, enabling the agent to incorporate predictive market signals. The goal is to learn a policy $\mathcal{\pi}:\mathit{S}\rightarrow\mathcal{A}$ that maximizes the expected discounted cumulative reward $\mathbb{E}[\sum^{\infty}_{t=0}\gamma^{t-1}r_{t}]$. At each trading day $t$ for specific stock $i$, the agent takes action $a_{i,t}:=\{a_{i,t}^{b},a_{i,t}^{h},a_{i,t}^{s}\}\in\mathcal{A}$ which means buy, hold and sell, according to the current environment $s_{t}\in\mathcal{S}$.
3.3. Vector Quantized Variational Autoencoder
VQ-VAE(van2017neural) is different from VAE (kingma2013auto) as it quantizes the observed data into a discrete token sequence. The VQ-VAE contains three parts, encoder, decoder, and discrete codebook, denoted by $\phi_{\text{enc}}$ with parameter $\theta_{\text{enc}}$, $\phi_{\text{dec}}$ with parameter $\theta_{\text{dec}}$ and $\mathcal{C}=\{(k,e_{k}\in R^{H})\}_{k=1}^{K}$ with parameter $\{e_{k}\in R^{H}\}_{k=1}^{K}$ respectively. Among them, $K$ is the size of the codebook, $H$ is the dimension of the vector, and $e_{k}$ is the $k$-th vector of the latent embedding space. Given an input $x$, the encoder $\phi_{\text{enc}}$ firstly encodes it into a set of continuous feature vectors $Z=\phi_{\text{enc}}(x)$ and quantization operation $q(\cdot)$ passes each feature vector $z\in Z$ through the discretization bottleneck following by selecting the nearest neighbor embedding in the codebook $\mathcal{C}$ as its discrete code sequence $Q_{z}$.
[…]
Then, the quantized feature for $z$ denoted as $Z^{q}_{z}$, is obtained through $Z^{q}_{z}=e_{Q_{z}}$. The quantized vectors $Z^{q}$ is fed into the decoder $\theta_{\text{dec}}$ to reconstruct the original data $x^{\prime}=\theta_{\text{dec}}(Z^{q})$.
Figure 2. The architecture of our proposed STORM model
4. Method
In this section, we represent our proposed latent factor model STORM, with the overall structure of the VQ-VAE depicted in Figure 2. In general, STORM models data into time-series and cross-sectional features, deriving factors by integrating both temporal and spatial perspectives. Firstly, we introduce the dual VQ-VAE architecture and explain the extraction of features in two modules. Then the factor module will fuse and align features and generate factor embeddings. Finally, we demonstrate the effectiveness of STORM through comprehensive evaluations on two critical downstream tasks: portfolio management and algorithmic trading.
4.1. Dual VQ-VAE Structure
The dual VQ-VAE architecture processes stock data from both spatial and temporal perspectives by patching the data within each module. The encoder then extracts time-series (TS) and cross-sectional (CS) features, which are subsequently clustered using a learnable codebook for quantized representation. Finally, the decoder reconstructs the data, effectively capturing complex patterns and enhancing model’s forecasting performance.
4.1.1. Patching
In stock data analysis, it is essential to extract features from both temporal and spatial dimensions. Time-series features capture the dynamics of a stock over time, while cross-sectional features reveal the correlations among different stocks. To achieve this, we partition data into patches (nie2022time) within the TS and CS modules. In the TS module, the observed data $\mathbf{x}$ is divided along the stock number dimension, with each patch containing data for one stock over $p$ days (as illustrated in Figure 2). Similarly, in the CS module, the observed data $\mathbf{x}$ is divided along the time axis, with each patch containing features of all stocks for a single trading day. Combining information from multi-scale features allows the model to capture both temporal and spatial dependencies, enhancing the depth and accuracy of market insights.
4.1.2. TS and CS Encoders
Considering the Transformer’s (vaswani2017attention) remarkable ability to capture both local and global dependencies via self-attention and patching strategies, it is particularly well-suited for long-term time-series forecasting (nie2022time). Therefore, we employ stacked transformer blocks as the encoders in both the TS and CS modules to capture complex time-series and cross-sectional patterns. This architecture enhances the accuracy and robustness of feature extraction, ultimately improving forecasting performance.
While transformers are widely applicable across domains such as natural language processing (NLP) and computer vision (CV), STORM’s flexible encoder-decoder framework allows for the integration of alternative architectures, such as long short-term memory (LSTM) (wang2021clvsa) and temporal convolutional networks (TCN) (yu2015multi).
An encoder–decoder configuration comprising 4 and 2 blocks, respectively, each with 4 and 8 attention heads, remains relatively lightweight, avoiding excessive computational cost and a substantial increase in model complexity.
4.1.3. Codebook Construction and Optimization
The construction of the codebook involves a learnable embedding space designed to quantize the encoded features into discrete representations (van2017neural). During training, the continuous latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$ are quantized into discrete space by mapping each vector to its nearest codebook entry, $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$, based on minimizing the Euclidean distance. This process translates the encoded features into discrete tokens, enabling the model to leverage a finite set of vectors to capture complex patterns.
The VQ-VAE offers significant advantages in the latent factor model, primarily through its use of discrete codebook embeddings: i) Discretization Benefits: The discretization process inherent in VQ-VAE helps in clustering distinct and meaningful factors, providing orthogonality and diversity which improves upon the factors derived from traditional methods like FactorVAE. ii) Explicit Factor Selection: The discrete token indices facilitate an explicit factor selection process, identifying the most relevant factors that influence stock returns and thus enhancing prediction accuracy. iii) Noise Reduction: By filtering out irrelevant or redundant information during factor selection, VQ-VAE mitigates noise, contributing to improved model robustness and more reliable financial insights.
To encourage equal usage of the codebook vectors (baevski2020wav2vec), we incorporate a diversity loss to ensure balanced representations across the codebook. It is achieved by maximizing the entropy of the averaged softmax distribution over vectors for each codebook $\bar{p}_{g}$:
(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}$
where $G$ represents the number of codebooks, set 2 in STORM. The diversity constraint helps enhance representational capacity, leading to better coverage of the input space.
However, the lack of independence among factors can introduce multicollinearity, making the model prone to overfitting and reducing its robustness across different market conditions. To address this issue, we impose an orthogonality constraint on the codebook, as recent studies (shin2023exploration) have shown that enforcing orthogonality allows discretized codes to maintain translation equivariance. The orthogonality loss is:
[…]
where $I_{k}\in\mathbb{R}^{K\times K}$ is the identity matrix, $\mathbf{e}$ is a surrogate for two codebook embeddings $\mathbf{ts}$ and $\mathbf{cs}$, $\ell_{2}(\mathbf{e})$ denotes $L_{2}$-normalized embeddings, and $\left|\left|\cdot\right|\right|_{F}$ is the Frobenius norm. The orthogonal constraint ensures that factors are independent of each other, allowing the effect of each factor on returns to be individually explained, which is crucial for asset pricing analysis.
4.2. TS and CS Decoders
The decoder is responsible for reconstructing the original data $\mathbf{x}$ from the quantized latent vectors $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$. We utilize Transformer as decoders either, to generate constructed data $\mathbf{x}_{ts}^{\prime}$ and $\mathbf{x}_{cs}^{\prime}$, which aim to closely approximate the original input data.
Finally, the encoders, decoders, and codebooks are jointly optimized by minimizing the following loss objectives:
[…]
where $sg[\cdot]$ is a stop-gradient operator.
4.3. Factor Module
The factor module aims to integrate the TS and CS latent features through feature fusion and alignment. Afterward, it utilizes prior-posterior learning to generate factor embeddings from latent features to predict stock future returns.
4.3.1. Feature Fusion and Alignment
As mentioned above we employ different patching methods to ensure that both time-series and cross-sectional features of the data were accurately captured. However, these methods result in two parallel processes that do not align at the feature level. To address the limitation, we employ a cross-attention mechanism to enhance the interaction and fusion of features at a fine-grained level, improving the model’s dynamic understanding of the input data. Additionally, contrastive learning is used at the semantic level to enhance feature representation by emphasizing semantic similarities and differences.
Cross-Attention Mechanism. In order to fuse the features in the TS and CS modules at a fine-grained level, we leverage a multi-scale encoder to combine information between two temporal and spatial patterns and produce stronger fusion features for factor generations, motivated by CrossViT (chen2021crossvit). The architecture incorporates cross-attention layers that explicitly learn correlations between two patterns, enabling the encoder to emphasize mutually informative regions while filtering out noise. This design balances accuracy and computational efficiency, ensuring the fused features yield richer and more discriminative signals for factor generation and that the resulting factors are more representative of market dynamics.
Contrastive Learning. To further enhance semantic consistency between TS and CS features, we integrate contrastive learning, drawing inspiration from its success in aligning multimodal representations (li2023blip; radford2021learning). A contrastive loss is introduced within the factor module to encourage features from TS and CS modules that represent similar market conditions to cluster closely in the embedding space, while pushing apart features corresponding to different conditions. This dual-stream alignment improves the coherence of the factor space and enhances the model’s ability to generalize across diverse market scenarios.
4.3.2. Prior-Posterior Learning
Through multi-scale encoder and contrastive learning layer, the CS and TS latent features are fused and aligned. In order to retain the codebook’s categorical properties for factors, we incorporate the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (dosovitskiy2020image), which are appended to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively. Then, two latent features concat together to get the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$.
Given the inherent volatility and complexity of the stock market, bridging the gap between noisy market data and an effective factor model remains a significant challenge (duan2022factorvae). Therefore, we introduce the prior-posterior structure to predict future returns and optimize latent factor embeddings. In the training stage, the posterior layer estimates posterior distribution of factor expected returns $\mathbf{z}_{post}$ from the true future stock returns $\mathbf{y}$ and the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$:
(7) $\mu_{post},\sigma_{post}]=\phi_{FE}(\mathbf{y},\mathbf{z}_{e}(\mathbf{x}))\text{,}$
and $\mathbf{z}_{post}$ follows the independent Gaussian distribution. In the inference stage, the prior layer only use latent factor embeddings without any future information leakage to calculate prior distribution of factor expected returns $\mathbf{z}_{prior}$:
(8) $\mu_{prior},\sigma_{prior}]=\phi_{FP}(\mathbf{z}_{e}(\mathbf{x}))\text{.}$
The return predictor computes future returns based on the factor expected return $\mathbf{z}$:
(9) $\hat{\mathbf{y}}=\bm{\alpha}+\sum_{k=1}^{K}\bm{\beta}^{k}\mathbf{z}^{k}+\bm{\epsilon}\text{.}$
The predicted returns $\hat{\mathbf{y}}$ can be directly integrated into portfolio optimization, algorithmic trading strategies, enabling dynamic rebalancing and improved risk management.
4.4. Downstream Tasks
Existing research on latent factor models (duan2022factorvae; wei2023hirevae) primarily focuses on predicting future returns by analyzing cross-sectional characteristics of market stocks, which means the profitability of the factor model in financial investment is typically evaluated through PM tasks. In contrast, STORM captures the cross-sectional features among stocks and the time-series features of individual stocks, enabling it to excel in both PM and single-asset AT tasks.
4.4.1. Portfolio Management
To evaluate the performance of the STORM and compare it with baseline methods, we follow the paradigm proposed by FactorVAE (duan2022factorvae), focusing on return prediction and the PM task. Specifically, we utilize the factor decoder network to generate stock future returns $\mathbf{\hat{y}}$, and then apply the TopK-Drop strategy [^3] https://qlib.readthedocs.io/en/latest/component/strategy.html, which constructs a daily portfolio by selecting the top $k$ stocks based on predicted returns, to backtest the factor model. Considering transaction costs in real markets, there is a turnover constraint that limits the number of changes in the portfolio composition between consecutive trading days. Specifically, when moving from $\mathbf{w}_{t}$ to $\mathbf{w}_{t+1}$, at most $d$ stocks can be replaced, where a replacement is defined as a stock whose weight changes from 0 (not held) to a positive value (newly bought) or from a positive value to 0 (fully sold). This constraint ensures that at least kdk-d stocks remain in the portfolio, reducing excessive turnover.
4.4.2. Algorithmic trading
In STORM, $\mathbf{z}_{e}(\mathbf{x})$ is the stock latent factor embeddings encoded by the dual VQ-VAE architecture, including both time-series and cross-sectional factor embeddings in the data. The latent factor embeddings $\mathbf{Z}$ are integrated into the observation set $\mathcal{O}=\{\mathbf{Z},\mathcal{R}\}$, where $\mathcal{R}$ is the reward function used to guide the agent’s learning and decision-making in the environment. By combining latent factor embeddings with reward, the agent can more accurately understand the market environment and adjust its strategy based on observations. In the AT task, since it involves trading a single stock, the features of other stocks are excluded. We use the Proximal Policy Optimization (PPO) algorithm (schulman2017proximal) to optimize the policy. In each iteration, the agent interacts with the environment to collect trajectories, computes advantage estimates, and updates the policy parameters via PPO’s clipped surrogate objective. This optimization process ensures the stability and efficiency of policy updates, helping the agent gradually improve the decision-making ability and achieve better trading strategies.
Table 1. Comparison of baselines across different tasks
Task Market ML & DL-based RL-based Factor model Ours
B&H LGBM (yang2020qlib) LSTM (yu2018forecasting) Transformer (yang2020qlib) SAC (haarnoja2018soft) PPO (schulman2017proximal) DQN (mnih2013playing) CAFactor (gu2021autoencoder) FactorVAE (duan2022factorvae) HireVAE (wei2023hirevae) STORM
Prediction
Portfolio Management
Algorithmic Trading
  • * indicates that the model is applicable, while signifies that it is NOT applicable to the corresponding task.
5. Experiment
In this section, we evaluate the proposed STORM on real stock markets and conduct extensive experiments to address the following research questions.
RQ1:
How does STORM perform on downstream tasks?
RQ2:
How to evaluate the effectiveness of STORM’s learned factors?
RQ3:
How do the key components contribute to the performance of STORM?
5.1. Experiment Settings
5.1.1. Datasets
We conduct experiments on two U.S. stock markets, SP500 and DJ30, using stock daily data that includes technical features based on Alpha158 (yang2020qlib). Both datasets span 16 years, from 2008-04-01 to 2024-03-31, encompassing global conditions, e.g., the 2007-2008 financial crisis and COVID-19. Datasets are chronologically divided into non-overlapping training (from 2008-04-01 to 2021-03-31) and test (from 2021-04-01 to 2024-03-31) sets.
5.1.2. Metrics
We compare STORM and baselines in terms of 6 financial metrics across the PM and AT tasks. The financial metrics include 2 profit criteria, annualized percentage yield (APY) and Cumulative Wealth (CW), 2 risk-adjusted profit criteria, Calmar ratio (CR) and annualized Sharpe ratio (ASR), and 2 risk criteria, maximum drawdown (MDD) and annualized volatility (AVO). The calculation formulas and meanings of these metrics are as follows:
  • CW is the total returns yielded from a portfolio strategy: $\text{CW}_{T}=\prod_{i=1}^{T}(1+r_{i})$, where $r_{i}$ is the net return.
  • APY measures the average wealth increment that one portfolio strategy could achieve compounded in a year, which is defined as $\text{APY}_{T}=\sqrt[y]{\text{CW}_{T}}-1$, where $y$ is the number of years corresponding to $T$ trading rounds.
  • MDD measures the largest loss from a historical peak in the cumulative wealth to show the worst case, which is defined as: $MDD=\mathop{\max}_{i=0}^{T}\frac{P_{i}-R_{i}}{P_{i}}$, where $R_{i}=\prod_{i=1}^{T}{\frac{V_{i}}{V_{i-1}}}$ and $P_{i}=\mathop{\max}_{i=1}^{T}R_{i}$.
  • AVO is the annualized standard deviation of daily returns and multiplied by $\sqrt{AT}$, where $AT$ is the average trading rounds of annual trading days and $AT=252$ for all the datasets.
  • ASR is an annualized volatility risk-adjusted return defined as $\text{ASR}=\frac{\text{APY}-R_{f}}{\text{AVO}}$, where $R_{f}$ is the risk-free return.
  • CR measures the drawdown risk-adjusted return of a portfolio calculated as $CR=\frac{\text{APY}}{\text{MDD}}$.
Typically, to evaluate the effectiveness of the learned factors in the PM task, we adopt the Rank Information Coefficient (RankIC) and the Information Ratio of RankIC (RankICIR):
  • RankIC is a ranking metric in finance, which measures the correlation between the predicted rankings and the actual returns. It is defined as:
    $\text{RankIC}_{s}=\frac{1}{N}\frac{(r_{\hat{y}_{s}}-\text{mean}(r_{\hat{y}_{s}}))^{T}(r_{y_{t}}-\text{mean}(r_{y_{s}}))}{\text{std}(r_{\hat{y}_{s}})\cdot\text{std}(r_{y_{s}})}$
    $\text{RankIC}=\frac{1}{T_{\text{test}}}\sum_{s=1}^{T_{\text{test}}}\text{RankIC}_{s}\text{,}$
    where $T_{test}$ denotes the number of trading days in the test range, $r_{y_{s}}$ and $r_{\hat{y}_{s}}$ represent the true and predicted ranks of stocks on the trading day $s$, respectively.
  • RankICIR is the information ratio of RankIC, which measures the stability of prediction,
    $\text{RankICIR}=\frac{\text{mean}(\text{RankIC}_{s})}{\text{std}(\text{RankIC}_{s})}\text{.}$
Table 2. Portfolio management task results of all models across six metrics (mean ± range, computed across 10 runs).
SP500 Dataset
Strategies Profit Risk-Adj. Profit Risk
[…]
Market Index 0.058 1.184 0.228 0.142 0.254 0.410
LightGBM 0.059 1.201 0.304 0.332 0.238 0.176
± 0.132 ± 0.456 ± 0.753 ± 0.755 ± 0.108 ± 0.008
LSTM 0.069 1.221 0.278 0.371 0.248 0.186
± 0.020 ± 0.068 ± 0.065 ± 0.114 ± 0.042 ± 0.013
Transformer 0.076 1.244 0.389 0.433 0.19811footnotemark: 1 0.174
± 0.028 ± 0.098 ± 0.165 ± 0.161 ± 0.057 ± 0.004
CAFactor 0.075 1.241 0.342 0.428 0.223 0.174
± 0.028 ± 0.100 ± 0.229 ± 0.165 ± 0.043 ± 0.006
FactorVAE 0.079 1.256 0.404 0.460 0.200 0.173
± 0.025 ± 0.085 ± 0.177 ± 0.145 ± 0.040 ± 0.007
HireVAE 0.077 1.249 0.361 0.448 0.216 0.172
± 0.029 ± 0.104 ± 0.180 ± 0.189 ± 0.048 ± 0.007
STORM 0.18822footnotemark: 2 1.683 1.189 1.052 0.166 0.171
± 0.055 ± 0.226 ± 0.661 ± 0.329 ± 0.050 ± 0.020
STORM-w/o-TS 0.090 1.300 0.503 0.570 0.181 0.175
± 0.074 ± 0.277 ± 0.390 ± 0.339 ± 0.043 ± 0.025
STORM-w/o-CS 0.089 1.294 0.623 0.592 0.146 0.167
± 0.045 ± 0.163 ± 0.385 ± 0.252 ± 0.073 ± 0.020
Improvement(%)33footnotemark: 3 137.97 34.00 194.31 128.70 26.26 2.91
DJ30 Dataset
Strategies Profit Risk-Adj. Profit Risk
[…]
Market Index 0.063 1.201 0.147 0.429 0.219 0.288
LightGBM 0.069 1.221 0.288 0.430 0.244 0.160
± 0.040 ± 0.140 ± 0.262 ± 0.267 ± 0.047 ± 0.007
LSTM 0.060 1.192 0.243 0.370 0.248 0.163
± 0.007 ± 0.023 ± 0.021 ± 0.043 ± 0.006 ± 0.005
Transformer 0.056 1.179 0.227 0.367 0.250 0.154
± 0.013 ± 0.043 ± 0.067 ± 0.085 ± 0.033 ± 0.002
CAFactor 0.059 1.186 0.233 0.382 0.252 0.153
± 0.015 ± 0.050 ± 0.058 ± 0.092 ± 0.021 ± 0.001
FactorVAE 0.076 1.246 0.352 0.480 0.225 0.159
± 0.038 ± 0.135 ± 0.349 ± 0.235 ± 0.063 ± 0.005
HireVAE 0.072 1.233 0.298 0.445 0.247 0.163
± 0.037 ± 0.132 ± 0.253 ± 0.256 ± 0.057 ± 0.012
STORM 0.148 1.517 1.396 1.052 0.108 0.140
± 0.046 ± 0.188 ± 0.679 ± 0.297 ± 0.026 ± 0.014
STORM-w/o-TS 0.079 1.259 0.621 0.603 0.127 0.142
± 0.057 ± 0.192 ± 0.425 ± 0.379 ± 0.038 ± 0.008
STORM-w/o-CS 0.073 1.236 0.533 0.566 0.138 0.140
± 0.051 ± 0.169 ± 0.400 ± 0.343 ± 0.034 ± 0.017
Improvement(%) 94.74 21.75 296.59 119.17 52.00 8.50
  • 11
    Underline indicates the best-performing baseline method result;
  • 22
    Bold indicates the best performance among our proposed models (including variants);
  • 33
    Improvement of STORM over the best-performing baselines.
Table 3. Algorithmic trading task results on all models across six standard metrics (averaged over 5 runs).
Models AAPL JPM IBM INTC MSFT
[…]
Buy&Hold 0.120 1.404 0.383 0.096 1.316 0.236 0.145 1.499 0.727 -0.117 0.690 -0.184 0.214 1.784 0.569
LightGBM 0.135 1.390 0.487 0.116 1.335 0.333 0.227 1.654 1.091 -0.042 0.880 0.038 0.267 2.068 0.637
LSTM 0.053 1.152 0.283 0.079 1.290 0.266 0.134 1.386 0.754 0.060 1.262 0.381 0.178 1.513 0.893
± 0.015 ± 0.042 ± 0.101 ± 0.019 ± 0.053 ± 0.023 ± 0.141 ± 0.406 ± 0.69 ± 0.119 ± 0.429 ± 0.62 ± 0.122 ± 0.352 ± 0.822
Transformer 0.083 1.240 0.512 0.133 1.384 0.614 0.131 1.377 0.782 0.079 1.290 0.458 0.138 1.397 0.726
± 0.099 ± 0.284 ± 0.864 ± 0.07 ± 0.201 ± 0.509 ± 0.072 ± 0.207 ± 0.541 ± 0.174 ± 0.68 ± 0.824 ± 0.118 ± 0.338 ± 0.643
DQN 0.135 1.374 0.510 0.105 1.305 0.607 0.139 1.400 0.802 0.061 1.185 0.442 0.166 1.475 0.534
± 0.075 ± 0.226 ± 0.239 ± 0.132 ± 0.383 ± 0.673 ± 0.055 ± 0.158 ± 0.270 ± 0.106 ± 0.305 ± 0.436 ± 0.035 ± 0.100 ± 0.107
SAC 0.147 1.509 0.528 0.131 1.383 0.400 0.207 1.598 1.170 0.056 1.165 0.353 0.229 1.656 0.929
± 0.021 ± 0.060 ± 0.052 ± 0.038 ± 0.111 ± 0.198 ± 0.057 ± 0.163 ± 0.335 ± 0.210 ± 0.6 ± 0.994 ± 0.114 ± 0.331 ± 0.585
PPO 0.137 1.379 0.496 0.128 1.372 0.356 0.146 1.422 0.779 -0.019 0.954 0.040 0.216 1.620 0.569
± 0.073 ± 0.209 ± 0.213 ± 0.026 ± 0.071 ± 0.059 ± 0.047 ± 0.136 ± 0.205 ± 0.177 ± 0.514 ± 0.418 ± 0.081 ± 0.232 ± 0.105
STORM 0.229 1.857 0.750 0.174 1.621 0.559 0.236 1.893 1.470 0.173 1.625 0.773 0.290 2.154 1.216
± 0.033 ± 0.154 ± 0.066 ± 0.032 ± 0.133 ± 0.081 ± 0.039 ± 0.184 ± 0.445 ± 0.067 ± 0.284 ± 0.293 ± 0.052 ± 0.262 ± 0.597
STORM-w/o-TS 0.199 1.730 0.766 0.127 1.437 0.627 0.154 1.536 1.294 0.106 1.212 0.267 0.254 1.979 0.964
± 0.059 ± 0.260 ± 0.274 ± 0.058 ± 0.214 ± 0.187 ± 0.023 ± 0.089 ± 1.012 ± 0.120 ± 0.145 ± 0.135 ± 0.060 ± 0.093 ± 0.503
STORM-w/o-CS 0.148 1.521 0.641 0.103 1.347 0.350 0.152 1.534 0.814 0.136 1.474 0.541 0.181 1.650 0.648
± 0.069 ± 0.283 ± 0.184 ± 0.043 ± 0.154 ± 0.106 ± 0.054 ± 0.219 ± 0.352 ± 0.082 ± 0.300 ± 0.328 ± 0.051 ± 0.221 ± 0.361
Improvement(%) 55.782 23.062 45.076 30.827 17.124 2.117 3.965 14.4501 20.408 118.987 26.969 66.594 8.614 4.159 30.893
Models AAPL JPM IBM INTC MSFT
[…]
Buy&Hold 0.447 0.313 0.269 0.405 0.406 0.237 0.679 0.199 0.213 -0.324 0.635 0.359 0.777 0.376 0.275
LightGBM 0.950 0.309 0.017 0.921 0.386 0.015 1.822 0.175 0.011 0.098 0.553 0.023 1.456 0.370 0.017
LSTM 0.623 0.244 0.012 0.801 0.335 0.012 1.268 0.163 0.010 0.287 0.149 0.010 1.345 0.246 0.014
± 0.163 ± 0.072 ± 0.003 ± 0.185 ± 0.061 ± 0.002 ± 0.924 ± 0.017 ± 0.001 ± 1.02 ± 0.101 ± 0.007 ± 0.842 ± 0.124 ± 0.003
Transformer 0.972 0.239 0.010 1.332 0.226 0.010 1.336 0.182 0.010 0.604 0.257 0.014 1.206 0.182 0.011
± 0.748 ± 0.187 ± 0.006 ± 0.607 ± 0.100 ± 0.003 ± 0.688 ± 0.100 ± 0.003 ± 0.789 ± 0.194 ± 0.006 ± 0.754 ± 0.134 ± 0.006
DQN 0.974 0.288 0.016 1.085 0.212 0.010 1.415 0.161 0.009 0.609 0.243 0.014 1.194 0.309 0.015
± 0.472 ± 0.159 ± 0.007 ± 1.270 ± 0.118 ± 0.003 ± 0.403 ± 0.049 ± 0.004 ± 0.584 ± 0.310 ± 0.009 ± 0.148 ± 0.079 ± 0.004
SAC 1.054 0.292 0.016 1.044 0.350 0.014 1.794 0.157 0.011 0.527 0.359 0.017 1.638 0.229 0.013
± 0.223 ± 0.022 ± 0.002 ± 0.276 ± 0.085 ± 0.001 ± 0.458 ± 0.068 ± 0.005 ± 1.061 ± 0.191 ± 0.004 ± 0.736 ± 0.215 ± 0.005
PPO 0.974 0.300 0.016 1.000 0.381 0.015 1.386 0.177 0.011 -0.044 0.511 0.019 1.276 0.353 0.017
± 0.412 ± 0.030 ± 0.001 ± 0.189 ± 0.018 ± 0.001 ± 0.231 ± 0.027 ± 0.003 ± 1.689 ± 0.115 ± 0.007 ± 0.299 ± 0.036 ± 0.001
STORM 1.346 0.282 0.016 1.245 0.299 0.014 2.107 0.137 0.009 1.189 0.227 0.015 1.672 0.210 0.015
± 0.162 ± 0.000 ± 0.001 ± 0.106 ± 0.016 ± 0.001 ± 0.302 ± 0.033 ± 0.001 ± 0.386 ± 0.055 ± 0.002 ± 0.201 ± 0.067 ± 0.002
STORM-w/o-TS 1.199 0.262 0.017 1.587 0.194 0.011 1.868 0.131 0.008 0.580 0.344 0.017 1.553 0.244 0.015
± 0.300 ± 0.052 ± 0.001 ± 0.296 ± 0.038 ± 0.001 ± 0.956 ± 0.077 ± 0.003 ± 0.229 ± 0.061 ± 0.000 ± 0.187 ± 0.093 ± 0.004
STORM-w/o-CS 1.027 0.229 0.015 0.858 0.327 0.013 1.397 0.177 0.010 0.914 0.312 0.018 1.401 0.295 0.016
± 0.326 ± 0.030 ± 0.001 ± 0.236 ± 0.064 ± 0.001 ± 0.405 ± 0.028 ± 0.001 ± 0.397 ± 0.136 ± 0.004 ± 0.224 ± 0.090 ± 0.002
Improvement(%) 27.704 4.184 - 19.144 8.491 - 15.642 16.561 11.111 95.238 - - 2.076 - -
5.1.3. Baselines
As shown in Table 1, we compare STORM with nine methods across one prediction task and two downstream tasks. The baseline methods fall into three categories: ML & DL-based models, Factor models, and RL-based models. We evaluate STORM against ML&DL models and factor models in prediction and PM tasks, and against ML&DL models and RL methods in the AT task. Moreover, STORM is a versatile all-rounder that supports three tasks, unlike other methods specialized in specific tasks.
  • Market
    • Buy-and-Hold (B&H) involves holding assets for an extended period, regardless of short-term market fluctuations, assuming that long-term returns will be more favorable.
  • ML&DL-based
    • LGBM (yang2020qlib) uses a series of tree models to predict price fluctuations and provide buy and sell signals.
    • LSTM (yu2018forecasting) utilizes long short-term memory to improve the accuracy of price predictions.
    • Transformer (yang2020qlib) models leverage self-attention mechanisms to enhance the precision of price forecasts.
  • RL-based
    • SAC (haarnoja2018soft) is an off-policy actor-critic algorithm using entropy regularization and soft value functions.
    • PPO (schulman2017proximal) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and efficiency.
    • DQN (mnih2013playing) approximates the action-value function and make trading decisions from market data with a deep Q-network.
  • Factor-model
    • CAFactor (gu2021autoencoder) learns latent factors from asset features.
    • FactorVAE (duan2022factorvae) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors.
    • HireVAE (wei2023hirevae) is a state-of-the-art VAE-based latent factor model capturing market and inter-stock factors.
5.1.4. Implementation Details
All experiments are implemented using PyTorch and conducted on two NVIDIA RTX H100 GPUs with 80GB of memory each. The input data is organized into batches with dimensions $(B,W,N,D)$, where $B=16$ is the batch size, $N$ is the number of stocks (28 for DJ30, 408 for SP500), $W=64$ denotes historical days, and $D=152$ is the total number of features, including the OHLC data and technical indicators.
The TS module uses patches of size $(4,1,D)$ (4-day sequences per stock), whereas the CS module uses $(1,N,D)$ (all stocks on a given day). Encoders (4 layers, 4 heads) and decoders (2 layers, 8 heads) are standard Transformer blocks with a shared embedding size of 256. Codebook sizes are searched in 128,256,512,1024,2048{128,256,512,1024,2048}, with 512 performing best.
[…]
For downstream tasks, we use Qlib’s [^4] https://github.com/microsoft/qlib TopK-Drop portfolio strategy (k=5k{=}5, d=3d{=}3) and PPO for trading. The pretrained TS/CS factors serve as states; all CS factors are retained while only TS factors relevant to traded stocks are used. The action space is buy, hold, sell. Rewards follow $r=\frac{v_{\text{post}}-v_{\text{pre}}}{v_{\text{pre}}}$ with $v=p\times m+c$, starting from $10^{6}$ initial cash and transaction cost $10^{-4}$.
The implementations of the ML & DL methods are based on Qlib (yang2020qlib). As for other baselines, we adopt the default settings in their public implementations. Each experiment is repeated multiple times with different random seeds, and the mean and variance of the metrics are reported.
5.2. Performance on Investment Tasks (for RQ1)
We use three types of metrics: profit, risk, and risk-adjusted profit to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading. The evaluation results are presented in Table 2 and Table 3.
For profitability, STORM outperforms all the baselines, with average improvements of 116.36% on APY over the best-performing baseline in the PM task and 43.64% over the best baseline methods on five stocks in the AT task. These results underscore the model’s capability to extract meaningful factors, and optimize decision-making for enhanced financial performance. For risk resistance, although STORM does not outperform the best baseline methods in terms of MDD and AVO on the INTC and MSFT datasets for the trading task, this can be attributed to the extreme price fluctuations during the COVID-19 event. Both stocks experienced a sharp decline of over 30%, followed by a subsequent rebound. Given STORM’s strong profitability capability, its trading strategy tends to aggressively capture trends, which naturally leads to larger drawdowns and increased volatility as a trade-off for higher potential returns.
Moreover, considering the return and risk simultaneously, STORM performs the best on ASR and CR with an average improvement of 123.94% and 245.45% respectively, over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task. Such an outstanding ability to balance return and risk is particularly valuable in real-world investment, where achieving high returns while effectively managing risk is crucial for long-term financial success.
Table 4. Factor quality evaluation task results on RankIC and RankICIR (mean ± range, computed across 10 runs).
Models SP500 DJ30
[…]
Future Return Prediction in Portfolio Management Task
LightGBM 0.027 0.274 0.031 0.272
± 0.006 ± 0.084 ± 0.005 ± 0.049
LSTM 0.034 0.333 0.031 0.329
± 0.006 ± 0.042 ± 0.004 ± 0.056
Transformer 0.035 0.340 0.033 0.343
± 0.007 ± 0.078 ± 0.005 ± 0.045
CAFactor 0.037 0.356 0.040 0.380
± 0.005 ± 0.084 ± 0.003 ± 0.043
FactorVAE 0.052 0.543 0.056 0.520
± 0.010 ± 0.122 ± 0.012 ± 0.081
HireVAE 0.057 0.558 0.058 0.563
± 0.006 ± 0.058 ± 0.006 ± 0.053
STORM 0.062 0.673 0.065 0.668
± 0.018 ± 0.155 ± 0.038 ± 0.287
STORM-w/o-TS 0.053 0.513 0.055 0.563
± 0.017 ± 0.145 ± 0.015 ± 0.127
STORM-w/o-CS 0.054 0.522 0.053 0.559
± 0.014 ± 0.118 ± 0.016 ± 0.156
Improvement(%)11footnotemark: 1 8.772 20.609 12.069 18.650
  • 11
    Improvement of STORM over the best-performing baselines.
5.3. Factor Quality (for RQ2)
To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task. The goal is not to pursue time-series forecasting accuracy, but to evaluate the predictive quality and robustness of the factors in capturing dynamic market patterns. In this task, STORM shows substantial improvements over six methods across the SP500 and DJ30 datasets. We use rank-based correlation metrics (i.e., RankIC, RankICIR) instead of error-based ones (e.g., MSE, MAE), as they better capture a factor’s ability to preserve the correct cross-sectional ordering of returns. As shown in Table  4, STORM achieves substantial average improvements of 14.690% on SP500 and 15.360% on DJ30, demonstrating its ability to capture more consistent and generalizable signals across diverse market conditions.
Figure 3. (a) Distribution of counts for codebook embedding indices. (b) Hyperparameter experiment results for different codebook sizes.
Factor Diversity. The diversity of factors is essential in financial investment, as it allows the model to capture the multi-dimensional characteristics of the market, avoiding redundancy or over-reliance on a single factor. In the codebook, each embedding represents a cluster center for a category of factor embeddings. To analyze the utilization rate of each factor category, we report the frequency of codebook embedding indices in the testing data. For clarity, we group every five indices into one interval, as illustrated in Figure 3(a). The results indicates a relatively even distribution of factor usage, highlighting the model’s effectiveness to maintain diversity during the factor selection process.
Codebook Design. Furthermore, we conduct a hyperparameter analysis on the DJ30 dataset to examine the effect of different codebook sizes, aiming to balance factor diversity and representational quality. An excessively large codebook may lead to sparse factor representations, limiting the effectiveness of factor utilization, while an overly small codebook could result in insufficient factor differentiation and information redundancy. Following standard practices, we evaluate five different codebook sizes, including 128, 256, 512, 1024, and 2048. As shown in Figure 3(b), a size of 512 achieves the best overall factor prediction performance, offering a favorable trade-off between expressiveness and efficiency.
5.4. Ablation Study (for RQ3)
We conducted ablation experiments with two simplified versions of STORM, namely STORM-w/o-TS and STORM-w/o-CS, which only extract cross-sectional and time-series factors, respectively. Specifically, the absence of either time-series or cross-sectional factors significantly reduces the model’s effectiveness across various tasks. In the stock return prediction task, STORM outperforms both simplified models, with an average improvement of 16.50% on the RankIC metric and 23.79% on the RankICIR metric, highlighting the enhanced accuracy achieved by incorporating both time-series and cross-sectional factors. In the PM and AT tasks, STORM and two simplified models all performed well. Notably, STORM-w/o-TS exhibited the lowest AVO among the three models, although its profitability did not surpass that of STORM. These findings suggest that methods like FactorVAE (duan2022factorvae), which focus solely on cross-sectional factors, may struggle to capture the full complexity of the financial market. Therefore, integrating both time-series and cross-sectional factors is essential for robust and accurate financial modeling.
5.5. Discussion on Performance
Efficiency. STORM runs efficiently in real-world daily trading without the ultra-low latency or specialized hardware needed for high-frequency systems, even on standard CPUs. This is enabled by a compact architecture with only a few transformer layers and 15.2M FP32 parameters on DJ30, enabling deployment on edge devices while supporting concurrent market analysis and trading.
Effectiveness. STORM delivers consistently strong performance across diverse tasks, including PM and AT, outperforming baselines that often specialize in one domain but fail to generalize. It balances return and risk, targeting comprehensive investment performance rather than over-optimizing for a single metric. Robustness and versatility across tasks justify the design of our proposed model.
Generalization. STORM generalizes across markets of varying scales and characteristics, excelling on both DJ30 (large-cap industrial blue chips) and SP500 (broad, diversified industries). These representative U.S. datasets cover concentrated and diversified indices, and their depth makes them a strong proxy for other markets like China or the U.K.. The flexible encoder-decoder architecture also allows using additional layers like Mamba  (gu2023mamba; mamba2) and ensures applicability to diverse trading environments and asset classes.
6. Conclusions and Future Work
In this paper, we propose STORM, a spatio-temporal latent factor model leveraging a dual VQ-VAE framework to learn both time-series and cross-sectional factors. STORM ensures factor orthogonality and diversity, crucial for effective factor selection in financial trading. Unlike traditional factors, STORM learns high-dimensional latent vectors that, though less interpretable, capture richer market structures. Extensive experiments demonstrate STORM’s superiority over the state-of-the-art methods across a prediction task and two downstream tasks, showcasing robust performance and effectiveness. In the future, we plan to extend STORM by incorporating side information like news, to explore the effects of exogenous factors.
Acknowledgments
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL). It is also supported, in part, by Jinan-NTU Green Technology Research Institute (GreenTRI).